A series RLC circuit has a capacitor with a capacitance of 48.0 μF , an inductor with an inductance of 0.800 H and a resistor with a resistance of 140. Ω. The circuit has a rms current of 6.90 A when the frequency is 56.0 Hz. What is the phase angle?
Answer #1

inductive reactance

XL = 2 pi f L = 2* 3.14* 56* 0.8 = 281.344 ohm

capacitive reactance

XC = 1/ ( 2 pi f C) = 1 / ( 2* 3.14* 56* 48* 10^-6) = 59.24 ohm

phase angle is given by

phi = arctan ( (XL - Xc) /R)

phi = arctan ( (281.344 - 59.24) / 140)

phi = 57.775

=========

Comment in case any doubt, will reply for sure. Goodluck